Consider a rectangle with perimeter 28 (units). Let the width of the 

rectangle be w (units) and let the Area of the region enclossed by the 

rectangle be A ( square units). 

 Express A as a function of w and state the domain and range of the 

 function.



1
Expert's answer
2021-05-30T23:45:38-0400

We know that the perimeter of rectangle can be given as,


"P = 2L+2W"


"L = \dfrac{P-2W}{2}"


"P = 28(units)"


Hence, "L = 14-W"


Now area "A" can be calculated as,


"A = L \times W"


"A = (14-W)W"


"A = 14W-W^2"


Domain : "0<W<14"


Range : "0<A<49"